### v2

# Geometric and Dynamic Scaling in Deep Transformers

Abstract
Scaling Transformer architectures to extreme depth often leads to rank collapse: representations become redundant and degenerate despite modern normalization schemes. We argue this is fundamentally a geometric problem. Standard residual connections implicitly assume monotonic feature accumulation is beneficial, but provide no mechanism to constrain update directions or erase outdated information. As depth increases, this causes uncontrolled drift from the semantic manifold and representational collapse.
We propose the Manifold-Geometric Transformer (MGT), a unified framework addressing these failures through two orthogonal principles. First, manifold-constrained hyper-connections (mHC) restrict residual updates to valid tangent space directions, preventing manifold drift. Second, deep delta learning (DDL) enables data-dependent, non-monotonic updates that support feature erasure rather than unconditional accumulation. Together, mHC controls update direction while DDL controls magnitude and sign, yielding stable geometric evolution across depth.
Our theoretical analysis predicts that coupling geometric constraints with dynamic erasure is essential for scaling beyond current depth limits. We design a rigorous evaluation protocol for ultra-deep networks ($100+$ layers) to test whether geometry, not depth itself, is the fundamental bottleneck in Transformer scalability.
1 Introduction
The residual connection is the cornerstone of modern Deep Learning, allowing gradients to flow through hundreds of layers. However, the standard additive update $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ makes a strong assumption: that the optimal update is always an unconstrained vector addition in Euclidean space. Recent geometric deep learning perspectives suggest that data flows on low-dimensional non-linear manifolds embedded in high-dimensional space. An unconstrained update $\mathcal{F}(\mathbf{x}_{l})$ often points “off-manifold,” leading to feature degradation and rank collapse in deep layers.
Simultaneously, the additive nature implies a “write-only” memory mechanism. It is optimizationally difficult for a network to “erase” information (i.e., learn $\mathcal{F}(\mathbf{x})\approx-\mathbf{x}$) when context shifts. This results in the “residual accumulation” problem, where noise accumulates over depth.
To resolve these dual challenges, we introduce the Manifold-Geometric Transformer (MGT). We propose that an ideal layer update should satisfy two conditions:
  1. 1.
    Geometric Validity (via mHC): The update vector should lie in the tangent space of the current manifold state ($T_{\mathbf{x}}\mathcal{M}$). This theoretically prevents the features from collapsing into degenerate subspaces.
  2. 2.
    Dynamic Traversal (via DDL): The step size along this tangent direction should be dynamic and reversible, allowing the model to perform “Erasure” (moving backwards) or “Identity” (staying still) operations explicitly.
Our contributions are formulated as follows:
  • We formulate Manifold-Constrained Hyper-Connections (mHC), a lightweight projection mechanism to regularize the output of Attention/FFN blocks.
  • We integrate Deep Delta Learning (DDL) with a dynamic gate $\beta$, enabling geometric Householder updates.
  • We provide a theoretical grounding for the synergy between mHC and DDL, arguing that they act as multipliers for signal integrity.
  • We propose a rigorous experimental framework designed to falsify the hypothesis that geometric constraints are essential for scaling Transformers beyond current depth limits.
2 Related Work
Rank Collapse in Deep Networks. dong2021attention provided seminal theoretical analysis showing that pure self-attention layers lose rank doubly exponentially with depth, fundamentally limiting representational capacity. This collapse occurs even with modern normalization techniques (ba2016layer; xiong2020layer), suggesting that architectural modifications beyond optimization tricks are necessary. Recent depth-scaling efforts (wang2022deepnet; Touvron_2021_ICCV) have proposed specialized initialization and normalization schemes to train networks exceeding 100 layers, yet these approaches treat symptoms rather than addressing the geometric root cause of collapse.
Residual Connections and Depth. Since ResNets (he2016deep), residual connections have enabled training of deep networks by providing gradient highways. However, the standard formulation $\mathbf{x}_{l+1}=\mathbf{x}_{l}+\mathcal{F}(\mathbf{x}_{l})$ implicitly assumes additive updates in Euclidean space are universally beneficial. As Touvron_2021_ICCV observed, this monotonic accumulation can lead to feature redundancy in very deep networks. Our work addresses this through explicit erasure mechanisms.
Manifold-Constrained Hyper-Connections (mHC). The recently proposed mHC framework (xie2025mhc) addresses stability in width-scaling by constraining hyper-connections to geometric manifolds (e.g., the Birkhoff polytope), preserving the identity mapping property of ResNets. We adopt this perspective as a “spatial regularizer” that projects updates onto the tangent space $T_{\mathbf{x}}\mathcal{M}$ of the data manifold. While xie2025mhc focused on connection geometry, we argue that dynamics along these connections require complementary mechanisms for feature redundancy.
Deep Delta Learning (DDL). Our dynamic update law builds on the recent Deep Delta Learning framework (zhang2026deep), which generalizes residual connections via a learnable Delta Operator—a rank-1 perturbation parameterized by reflection direction $k(\mathbf{x})$ and gate $\beta(\mathbf{x})$. The key insight is that $\beta$ enables interpolation between identity ($\beta\to 0$), projection ($\beta\to 1$), and reflection ($\beta\to 2$), providing an explicit mechanism for information erasure that allows networks to actively forget outdated features.
Synergizing Geometry and Dynamics. While DDL provides erasure capability, applying it in unconstrained Euclidean space risks optimization difficulties. Conversely, mHC provides manifold constraints but lacks backward traversal (erasure) along the manifold. Our MGT architecture unifies these orthogonal contributions: mHC defines valid directions while DDL controls magnitude and sign, achieving stable “Erasure-and-Write” dynamics for ultra-deep scaling.
3 Methodology
We formulate the proposed Manifold-Geometric Transformer (MGT) block. We begin by establishing the geometric premise of feature propagation in deep networks, followed by the detailed derivation of our two core components.
3.1 Architecture Overview
We redesign the fundamental building block of the Transformer to explicitly separate feature generation from feature propagation. As illustrated in Figure 1, the proposed Manifold-Geometric Transformer (MGT) Block deviates from the standard Post-LN structure by introducing a geometric processing stage before the residual addition.
The forward pass of an MGT layer $l$ is decomposed into three distinct phases:
  1. 1.
    Raw Feature Generation: First, the input state $\mathbf{X}_{l}$ is normalized and processed by a standard mixing module (Multi-Head Self-Attention or Feed-Forward Network) to generate a candidate update vector $\mathbf{V}_{raw}$.
  2. 2.
    Geometric Rectification (via mHC): Instead of direct addition, $\mathbf{V}_{raw}$ is passed through the Manifold-Constrained Hyper-Connection (mHC) module. This module acts as a spatial filter, projecting the raw update onto the estimated tangent space of the data manifold.
  3. 3.
    Delta Dynamics (via DDL): Finally, the Deep Delta Learning (DDL) controller computes a dynamic gating scalar $\bm{\beta}$ based on the input context $\mathbf{X}_{l}$. This scalar modulates the rectified vector. Crucially, to strictly approximate a geometric reflection, we include a subtraction term proportional to the current state projection, allowing for true “erasure” mechanics.
[…]
Figure 1: Architecture of the Manifold-Geometric Transformer (MGT) Block. The pipeline explicitly separates (1) feature generation, (2) geometric rectification (blue/purple), and (3) dynamic erasure (orange). The dashed orange line illustrates the context-aware gating mechanism derived directly from the input state.
3.2 Preliminaries: The Manifold Hypothesis in Deep Layers
Let $\mathcal{X}=\{\mathbf{x}_{l}\}_{l=0}^{L}$ denote the sequence of hidden states in a Transformer, where $\mathbf{x}_{l}\in\mathbb{R}^{S\times D}$. We posit that valid semantic representations do not populate the entire Euclidean space $\mathbb{R}^{D}$, but rather lie on a lower-dimensional Riemannian manifold $\mathcal{M}\subset\mathbb{R}^{D}$. A naive step $\mathbf{x}_{l}+\mathbf{v}_{l}$ often forces the representation off the manifold ($\mathbf{x}_{l+1}\notin\mathcal{M}$). To maintain structural integrity, the update vector must be constrained to the local tangent space $T_{\mathbf{x}_{l}}\mathcal{M}$.
3.3 Manifold-Constrained Hyper-Connections (mHC)
The goal of mHC is to rectify the raw update vector $\mathbf{v}_{raw}$ generated by the sub-layers (MHSA or FFN). Since computing the exact tangent space via eigendecomposition is computationally prohibitive, we employ an efficient Soft Subspace Approximation.
We define the mHC operator $\Psi(\cdot)$ as a learnable gating mechanism that approximates orthogonal projection:
$\mathbf{v}_{mHC}=\mathbf{v}_{raw}\odot\sigma(\text{LN}(\mathbf{W}_{gate}\mathbf{x}_{l}))$ (1)
where $\mathbf{W}_{gate}$ projects the input to a semantic importance score. While not a strict orthogonal projection in the linear algebra sense, this acts as a soft manifold constraint, suppressing feature dimensions (noise subspace) that deviate from the current semantic trajectory defined by $\mathbf{x}_{l}$.
3.4 Deep Delta Learning (DDL): The Geometric Update Law
Once the update direction is rectified by mHC, we must determine the update dynamics. Standard connections assume a fixed step size. To enable true erasure (moving backwards along the current state vector), we adopt the generalized Householder form (zhang2026deep).
We introduce the Delta Controller $\beta(\mathbf{x}_{l})$:
$\bm{\beta}=\lambda\cdot\tanh(\mathcal{G}(\mathbf{x}_{l}))+\epsilon$ (2)
where $\mathcal{G}(\cdot)$ is a lightweight linear projection.
The Generalized Householder Update: To strictly implement the geometric capability of “erasing” the current state information, the update rule must include a subtractive term proportional to the current state. We formulate the final update as:
$\mathbf{x}_{l+1}=\mathbf{x}_{l}+\bm{\beta}\odot(\mathbf{v}_{mHC}-\alpha\cdot\text{Proj}_{\mathbf{x}}(\mathbf{x}_{l}))$ (3)
In our simplified implementation, we approximate the projection term $\text{Proj}_{\mathbf{x}}(\mathbf{x}_{l})$ with the normalized state itself, governed by a learnable scalar $\alpha$. This formulation allows:
  • Accumulation: When $\beta>0$ and $\mathbf{v}_{mHC}$ dominates.
  • Erasure/Reflection: When $\beta$ scales the negative term $\alpha\mathbf{x}_{l}$, effectively subtracting current features from the residual stream (Householder-like reflection).
3.5 Algorithm
Algorithm 1 summarizes the forward pass of a single MGT block. The procedure explicitly separates the four computational phases: (1) standard feature extraction via LayerNorm and the mixing sub-layer, (2) geometric rectification through the mHC projection, (3) dynamic gate computation via the DDL controller, and (4) the final Householder-style update that combines the rectified features with the erasure term. This modular decomposition enables independent analysis of each component’s contribution and facilitates efficient ablation studies.
Algorithm 1 Forward Pass of Manifold-Geometric Transformer Block
0: Input state $\mathbf{X}_{l}\in\mathbb{R}^{S\times D}$, Sub-layer $\mathcal{F}$
0: Updated state $\mathbf{X}_{l+1}$
1:1. Raw Generation
[…]
4:2. mHC Projection (Geometry)
[…]
7:3. DDL Dynamics (Kinematics)
[…]
10:4. Geometric Update (Householder)
[…]
13:return $\mathbf{X}_{l+1}$
4 Proposed Evaluation Framework
We design a comprehensive experimental protocol consisting of five experiments to rigorously validate the Manifold-Geometric Transformer (MGT). The goal is not merely to achieve state-of-the-art accuracy, but to falsify our core hypothesis: that constraining updates to the tangent space while enabling erasure dynamics leads to superior signal propagation in deep networks.
4.1 Experimental Setup
Datasets. We evaluate on language modeling benchmarks:
  • WikiText-2/103: Standard language modeling benchmarks for measuring perplexity and training dynamics.
  • Synthetic Copy Task: A controlled task where the model must copy the first half of the sequence to the second half, enabling precise analysis of information flow.
Baselines. To isolate the contribution of MGT components:
  • Standard: Post-LN Transformer as the control group.
  • +mHC Only: Standard Transformer augmented with mHC projection.
  • +DDL Only: Standard Transformer augmented with DDL dynamics.
  • MGT (Full): Complete architecture with both mHC and DDL.
4.2 Experiment 1: Rank Evolution Analysis
This experiment validates our core hypothesis regarding representational collapse in ultra-deep networks.
Setup: We compare Standard Transformer and MGT across depths $L\in\{12,24,48,100\}$ with model dimension $d=256$. Crucially, we measure effective rank on trained models (not random initialization) after 50 epochs on WikiText-2, as rank collapse manifests during optimization.
Metric: Effective Rank. We compute the normalized effective rank of the hidden state matrix $\mathbf{X}_{l}\in\mathbb{R}^{S\times D}$ at layer $l$:
$\text{Rank}_{eff}(\mathbf{X}_{l})=\frac{\exp(H(\bm{\sigma}(\mathbf{X}_{l})))}{S}$ (4)
where $H(\bm{\sigma})=-\sum_{i}\hat{\sigma}_{i}\log\hat{\sigma}_{i}$ is the Shannon entropy of the normalized singular values $\hat{\sigma}_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$.
Hypotheses:
  • H1: Standard Transformer exhibits monotonic rank decay, with $\text{Rank}_{eff}\to 0$ as $L\to 100$.
  • H2: MGT maintains $\text{Rank}_{eff}>0.5$ even at 100 layers.
Derived Metrics: We compute the Rank Preservation Ratio $\rho=\text{Rank}_{eff}(L)/\text{Rank}_{eff}(0)$ and the Rank Decay Rate as the slope of log-rank vs. depth. Results are averaged over 3 seeds.
4.3 Experiment 2: Ablation Study (Synergy Verification)
This experiment quantifies the individual and combined contributions of mHC and DDL.
Setup: We train four model variants with $L=48$ layers on WikiText-2 for 50 epochs with identical hyperparameters ($d=256$, batch size 64, learning rate $10^{-3}$). We use 48 layers to ensure rank collapse is observable in the baseline. Results are averaged over 3 seeds (Table 1).
Table 1: Ablation Protocol. Four configurations isolate the geometric and dynamic contributions.
Configuration mHC DDL Expected Behavior
1. Standard - - Control group (baseline loss).
2. +mHC Only - Improved stability, limited erasure.
3. +DDL Only - Erasure capability, potential drift.
4. MGT (Full) Optimal: stable geometry + dynamic erasure.
Synergy Coefficient. We define $\mathcal{S}$ to measure super-additive effects:
$\mathcal{S}=\underbrace{(\mathcal{L}_{base}-\mathcal{L}_{MGT})}_{\text{MGT Gain}}-\underbrace{(\mathcal{L}_{base}-\mathcal{L}_{mHC})}_{\text{mHC Gain}}-\underbrace{(\mathcal{L}_{base}-\mathcal{L}_{DDL})}_{\text{DDL Gain}}$ (5)
A positive $\mathcal{S}>0$ validates the orthogonal synergy hypothesis.
4.4 Experiment 3: Beta Distribution Analysis
This experiment validates that DDL learns meaningful erasure dynamics.
Setup: We train a 100-layer MGT on WikiText-2 for 100 epochs and analyze the distribution of learned $\bm{\beta}$ values at epochs $\{0,25,50,100\}$. The deep architecture ensures that erasure dynamics have sufficient opportunity to emerge.
Hypotheses (Phase Transition):
  • Early Layers: $\mathbb{E}[\beta]>0$ (accumulation mode for low-level features).
  • Deep Layers: Significant fraction of $\beta<0$ (erasure mode for semantic refinement).
  • Training Dynamics: The transition point should sharpen as training progresses.
Metrics: Per-layer statistics including $\mathbb{E}[\beta]$, $\text{Var}[\beta]$, and the fraction of negative gates $P(\beta<0)$.
4.5 Experiment 4: Depth Scaling
This experiment tests whether MGT scales more favorably with depth than Standard Transformers.
Setup: To ensure fair comparison, we fix the total parameter count at ${\sim}20$M and vary depth $L\in\{24,48,100,200\}$, adjusting width accordingly. Each configuration is trained for 50 epochs on WikiText-2 with 3 seeds. We use gradient checkpointing to enable training of 200-layer models on a single A100.
Metrics:
  • Final Perplexity: Lower is better.
  • Convergence Speed: Epochs to reach target perplexity.
  • Training Stability: Variance of loss across seeds.
Hypothesis: MGT should exhibit favorable depth scaling, achieving lower perplexity at greater depths where Standard Transformers degrade.
4.6 Experiment 5: Language Modeling
This experiment provides end-to-end validation on a realistic benchmark.
Setup: We train both Standard Transformer and MGT on WikiText-103 for 100 epochs with matched architectures ($d=512$, $L=24$, $h=8$, ${\sim}50$M parameters). We additionally evaluate on the larger OpenWebText corpus to test scaling behavior.
Metrics:
  • Perplexity (PPL): Primary evaluation metric on validation set.
  • Learning Curves: Training and validation loss trajectories.
  • Parameter Overhead: mHC adds ${\sim}2L\cdot d^{2}$ parameters; DDL adds ${\sim}2L\cdot d^{2}$ parameters (total ${\sim}25\%$ overhead).
Expected Outcome: MGT should achieve lower perplexity with improved training stability. The gap should widen with increased depth.
5 Conclusion
In this work, we introduced the Manifold-Geometric Transformer (MGT), a theoretical framework that unifies two disparate concepts in deep learning: manifold constraints and residual dynamics. By mathematically formalizing the residual update as a navigation problem on a latent manifold, we derived the need for orthogonal projection (mHC) combined with signed geometric updates (DDL).
While empirical validation is the immediate next step, our theoretical analysis and proposed experimental design suggest that MGT offers a principled solution to the long-standing problems of rank collapse and residual accumulation. We believe this geometric perspective—viewing layer updates as controlled vectors in a tangent bundle—opens new avenues for designing ultra-deep, mathematically robust neural architectures.